OpenStudy (anonymous):

The centrifugal force acting on a mass traveling in a circular path varies directly as the square of its speed and inversely as the radius of its circle. If a 1200 pound force is acting on a mass traveling 40 feet per second along a path that is 6 feet in diameter, what speed in feet per second will the same mass travel on the same path with a 90 pound force acting on it? Please show work!

jimthompson5910 (jim_thompson5910):

Step 1) Translation How would you translate "The centrifugal force acting on a mass traveling in a circular path varies directly as the square of its speed and inversely as the radius of its circle"

OpenStudy (anonymous):

y = kx^2 y = k/x?

jimthompson5910 (jim_thompson5910):

good, but I should backtrack a bit

jimthompson5910 (jim_thompson5910):

Step 1) Identity the variables Let F = force (or centrifugal force) s = speed r = radius

jimthompson5910 (jim_thompson5910):

then step 2) Translation ""The centrifugal force acting on a mass traveling in a circular path varies directly as the square of its speed" ----> F = ks^2 and "The centrifugal force acting on a mass traveling in a circular path varies... inversely as the radius of its circle" ---> F = k/r So combine the two to get F = (ks^2)/r

OpenStudy (anonymous):

Wait, wouldn't it be k(ks^2)/r? Since it's k/r? I'm not quite completely following when combining them.

jimthompson5910 (jim_thompson5910):

basically we have F = ks^2 and F = k/r and we "combine" the ks into one, but we're kinda abusing notation because the k's are different So we should have F = ms^2 and F = n/r where m and n are constants Combine the two: F = (ms^2)*(n/r) = (m*n*s^2)/r = (ks^2)/r Where k is a constant and k =m*n So even though the prev line of reasoning was off, it still gives us the same equation.

OpenStudy (anonymous):

Oh. ok. I get what happened now. :)

jimthompson5910 (jim_thompson5910):

alright, that's great

jimthompson5910 (jim_thompson5910):

So do you know where to go from here?

OpenStudy (anonymous):

Do I just plug in numbers? 1200 = mn40^2/3 3600 = 1600mn mn = 9/4 90 = 9/4(s^2)/3 270 = 9/4(s^2) 120 = s^2 s = 2√30?

jimthompson5910 (jim_thompson5910):

You can use k to make things simpler, but it will give you the same answer.

jimthompson5910 (jim_thompson5910):

In either case, you get k = 9/4 (or mn = 9/4) and s = 2*sqrt(30), so very nice job

OpenStudy (anonymous):

Thank you! :)

jimthompson5910 (jim_thompson5910):

Also, you might want to convert everything to decimal, so 2*sqrt(30) = 10.954451 roughly

OpenStudy (anonymous):

Oh. I'm doing this for a math team question and we're not allowed to use calculators which is why I left it in simplest radical form.

jimthompson5910 (jim_thompson5910):

oh gotcha, then you got it